OpenStudy (anonymous):

What law states that when there is an interaction between charged objects there is a non-contact force which acts over some distance of separation? A. Faraday's law B. Ohm's law C. Coulomb's law D. Interaction law

OpenStudy (unklerhaukus):

\[|F_{q_1q_2}|=\frac 1{4\pi \epsilon_0}\frac{|q_1q_2|}{r^2} \]

OpenStudy (anonymous):

so wat is tht?

OpenStudy (anonymous):

that is coulomb's law

OpenStudy (anonymous):

oh okay thnk u @vasudev

OpenStudy (anonymous):

c. Coulomb's Law